### v1

# Armadillo and Eigen: A Tale of Two Linear Algebra Libraries

[…]
mm[#1]
Armadillo and Eigen: A Tale of Two Linear Algebra Libraries
Mauricio Vargas Sepúlveda (ORCID $0$-0003-1017-7574)
Department of Political Science, University of Toronto
Munk School of Global Affairs and Public Policy, University of Toronto
Corresponding author: m.sepulveda@mail.utoronto.ca
[…]
1 Abstract
This article introduces cpp11eigen, a new R package that integrates the powerful Eigen C++ library for linear algebra into the R programming environment. This article provides a detailed comparison between Armadillo and Eigen speed and syntax. The cpp11eigen package simplifies a part of the process of using C++ within R by offering additional ease of integration for those who require high-performance linear algebra operations in their R workflows. This work aims to discuss the tradeoff between computational efficiency and accessibility.
2 Introduction
R is widely used by non-programmers (Wickham et al. 2019), and this article aims to introduce benchmarks in a non-technical yet formal manner for social scientists. Our goal is to provide a fair comparison between Eigen and Armadillo, being both highly efficient linear algebra libraries written in C++. We do it by using cpp11eigen and cpp11armadillo.
Eigen emphasizes flexibility and speed, while Armadillo focuses on a balance between speed and easy of use.
‘RcppEigen’, introduced in 2011, integrates Eigen with R through the Rcpp package, enabling the use of C++ for performance-critical parts of R code. ‘RcppArmadillo’ has a similar goal (Sanderson and Curtin 2016; Eddelbuettel and Sanderson 2014). At the time of writing this article, 247 CRAN packages depend on ‘RcppEigen’, 755 on ‘RcppArmadillo’ (Lee 2024), and therefore these are highly successful packages.
cpp11eigen is an independent project that aims to simplify the integration of R and C++ by using ‘cpp11’, an R package that eases using C++ functions from R. A distinctive characteristics of cpp11eigen is the vendoring capability, meaning that it allows to copy its code into a project, making it a one-time dependency with a fixed and stable code until it is updated, and it is useful in restricted environments such as servers and clusters (Wickham et al. 2019; Vaughan, Hester, and François 2023).
cpp11armadillo offers similar features and both libraries are useful in cases where vectorization (e.g., applying an operation to a vector or matrix as a whole instead of looping over each element) is not possible or challenging. A detailed discussion and examples about why and when (and when not) rewriting R code in C++ is useful can be found in Burns (2011) and Vargas Sepúlveda (2023).
We followed four design principles when developing cpp11eigen, same as cpp11armadillo (Vargas Sepúlveda and Schneider Malamud 2024): column oriented, package oriented, header-only, and vendoring capable.
3 Syntax and benchmarks
One possibility is to start by creating minimal R packages with the provided templates.
remotes::install_github("pachadotdev/cpp11armadillo")
remotes::install_github("pachadotdev/cpp11eigen")
cpp11eigen::create_package("armadillobenchmark")
cpp11eigen::create_package("eigenbenchmark")
Comparing numerical libraries requires to write equivalent codes. For instance, in R we can use apply() functions while in C++ we need to write a for loop, and this allows a fair comparison between the two libraries. However, R has heavily optimized functions that also verify the input data, such as lm() and glm(), that do not have a direct equivalent in Armadillo or Eigen, and for a fair comparison the options are to write a simplified function for the linear model in R or to write a more complex function in C++.
The ATT benchmark, is a set of functions that can be rewritten using Armadillo and Eigen with relative ease, and test has the advantage of being well-known and widely used in the R community.
The first test in the ATT benchmark is the creation, transposition and deformation of an $N\times N$ matrix ($500\times 2,$ 500 × 2 , 500 in the original test). The R code comparable to the Armadillo code is:
matrix_calculation_01_r <- function(n) {
  a <- matrix(rnorm(n * n) / 10, ncol = n, nrow = n)
  b <- t(a)
  dim(b) <- c(n / 2, n * 2)
  a <- t(b)
  return(0L)
The Armadillo code is very similar:
#include <cpp11.hpp>
#include <cpp11armadillo.hpp>
using namespace arma;
using namespace cpp11;
[[cpp11::register]] int matrix_calculation_01_arma_(const int& n) {
  mat a = randn<mat>(n,n) / 10;
  mat b = a.t();
  b.reshape(n/2, n*2);
  a = b.t();
  return 0;
The Eigen code requires to create a function to draw random numbers from a normal distribution but it has a built-in function for the uniform distribution:
#include <cpp11.hpp>
#include <cpp11eigen.hpp>
#include <random>
using namespace Eigen;
using namespace cpp11;
std::mt19937& random_normal() {
  static std::random_device rd;
  static std::mt19937 gen(rd());
  return gen;
[[cpp11::register]] int matrix_calculation_01_eigen_(const int& n) {
  std::normal_distribution<double> d(0, 1);
  MatrixXd a = MatrixXd::NullaryExpr(n, n, [&]() {
    return d(random_normal());
  }) / 10;
  // for the uniform distribution this is simpler
  // MatrixXd a = MatrixXd::Random(n, n) / 10;
  MatrixXd b = a.transpose();
  b.resize(n / 2, n * 2);
  return 0;
The functions to do not move data between R and C++, and this is intentional to focus on the performance of the linear algebra libraries and not adding overhead from data transfer in the benchmarks. Each function creates a matrix and conducts equivalent operations on it. The returned value is zero in R and C++ in case that the functions run without errors.
The benchmarks were conducted on a ThinkPad X1 Carbon Gen 9 with the following specifications:
  • Processor: Intel Core i7-1185G7 with eight cores
  • Memory: 16 GB LPDDR4Xx-4266
  • Operating System: Pop!_OS 22.04 based on Ubuntu 22.04
  • R Version: 4.4.1
  • BLAS Library: OpenBLAS 0.3.20
The median times for the adapted and comparable implementations of the ATT benchmarks are as follows:
Table 1: Matrix calculation
Operation Time (s) Rank
$400\times 2,$ 400 × 2 , 400 $\text{matrix}^{1,000}$ Armadillo $0.188$ 1
$400\times 2,$ 400 × 2 , 400 $\text{matrix}^{1,000}$ Eigen $0.301$ 2
$400\times 2,$ 400 × 2 , 400 $\text{matrix}^{1,000}$ R $0.325$ 3
$800\times 2,$ 800 × 2 , 800 cross-product matrix - Armadillo $0.398$ 1
$800\times 2,$ 800 × 2 , 800 cross-product matrix - R $0.444$ 2
$800\times 2,$ 800 × 2 , 800 cross-product matrix - Eigen $1.151$ 3
Creation and modification of a $500\times 2,$ 500 × 2 , 500 matrix - Armadillo $0.204$ 1
Creation and modification of a $500\times 2,$ 500 × 2 , 500 matrix - Eigen $0.232$ 2
Creation and modification of a $500\times 2,$ 500 × 2 , 500 matrix - R $0.294$ 3
Linear regression over a $000\times 3,$ 000 × 3 , 000 matrix - Armadillo $0.459$ 1
Linear regression over a $000\times 3,$ 000 × 3 , 000 matrix - R $5.303$ 2
Linear regression over a $000\times 3,$ 000 × 3 , 000 matrix - Eigen $8.809$ 3
Sorting of 7,000,000 values - Armadillo $0.663$ 1
Sorting of 7,000,000 values - Eigen $0.691$ 2
Sorting of 7,000,000 values - R $0.759$ 3
Table 2: Matrix functions
Operation Time (s) Rank
Cholesky decomposition of a $000\times 3,$ 000 × 3 , 000 matrix - Armadillo $0.608$ 1
Cholesky decomposition of a $000\times 3,$ 000 × 3 , 000 matrix - R $0.709$ 2
Cholesky decomposition of a $000\times 3,$ 000 × 3 , 000 matrix - Eigen $2.902$ 3
Determinant of a $500\times 2,$ 500 × 2 , 500 matrix - Armadillo $0.293$ 1
Determinant of a $500\times 2,$ 500 × 2 , 500 matrix - R $0.303$ 2
Determinant of a $500\times 2,$ 500 × 2 , 500 matrix - Eigen $0.562$ 3
Eigenvalues of a $640\times 640$ matrix - Armadillo $0.367$ 1
Eigenvalues of a $640\times 640$ matrix - R $0.369$ 2
Eigenvalues of a $640\times 640$ matrix - Eigen $1.629$ 3
Fast Fourier Transform over 2,400,000 values - Eigen $0.14$ 1
Fast Fourier Transform over 2,400,000 values - R $0.23$ 2
Fast Fourier Transform over 2,400,000 values - Armadillo $0.294$ 3
Inverse of a $600\times 1,$ 600 × 1 , 600 matrix - Armadillo $0.312$ 1
Inverse of a $600\times 1,$ 600 × 1 , 600 matrix - R $0.324$ 2
Inverse of a $600\times 1,$ 600 × 1 , 600 matrix - Eigen $0.758$ 3
Table 3: Programmation
Operation Time (s) Rank
3,500,000 Fibonacci numbers calculation - Eigen $1.4\times 10^{-1}$ 1
3,500,000 Fibonacci numbers calculation - Armadillo $1.7\times 10^{-1}$ 2
3,500,000 Fibonacci numbers calculation - R $1.7\times 10^{-1}$ 3
Creation of a $000\times 3,$ 000 × 3 , 000 Hilbert matrix - Eigen $4.6\times 10^{-6}$ 1
Creation of a $000\times 3,$ 000 × 3 , 000 Hilbert matrix - Armadillo $5.9\times 10^{-2}$ 2
Creation of a $000\times 3,$ 000 × 3 , 000 Hilbert matrix - R $1.5\times 10^{-1}$ 3
Creation of a $500\times 500$ Toeplitz matrix - Eigen $7.9\times 10^{-7}$ 1
Creation of a $500\times 500$ Toeplitz matrix - Armadillo $4\times 10^{-4}$ 2
Creation of a $500\times 500$ Toeplitz matrix - R $2.6\times 10^{-3}$ 3
Escoufier’s method on a $45\times 45$ matrix - Armadillo $2.4\times 10^{-2}$ 1
Escoufier’s method on a $45\times 45$ matrix - Eigen $3.2\times 10^{-2}$ 2
Escoufier’s method on a $45\times 45$ matrix - R $1.4\times 10^{-1}$ 3
Grand common divisors of 400,000 pairs - Eigen $2.1\times 10^{-2}$ 1
Grand common divisors of 400,000 pairs - Armadillo $2.3\times 10^{-2}$ 2
Grand common divisors of 400,000 pairs - R $1.884$ 3
The results reveal that Armadillo leads in most of the benchmarks, but Eigen is particularly faster in some tests such as the Fast Fourier Transform. R is the second or third in all benchmarks, but it is important to note that R comes with an additional advantage in terms of simplified syntax and the ability to run the code without compiling it.
These tests are not exhaustive, and we must be cautious when interpreting the results. The ATT benchmark is a good starting point, but it does not cover mundane tasks such as data manipulation, and it is important to consider the tradeoff between computational efficiency and ease of use.
4 Comparison with R packages
The syntax and speed differences posit a similar case to the tradeoff between using dplyr and data.table (Wickham et al. 2019; Barrett et al. 2024), where dplyr is easier to use but data.table is faster. dplyr was not designed to be fast but data.table was not designed to be easy to use. For instance, the code to obtain the grouped means by number of cylinders in the mtcars dataset is:
# dplyr
mtcars %>%
  group_by(cyl) %>%
  summarise_all(mean)
# data.table
as.data.table(mtcars)[, lapply(.SD, mean), by = cyl]
The benchmark for the grouped means reveals that dplyr has a median time of 2.7 ms and data.table has a median time of 600 $\mu$s, and this means that dplyr is four times slower than data.table at this task. The syntax of dplyr is easier to understand for non-programmers, but data.table can be equally expressive for users who are familiar with its syntax.
The tests for Armadillo and Eigen reveal that, for repeated and computationally intensive tasks, rewriting R code in C++ can lead to significant performance improvements, but it comes at the cost of learning a new syntax.
As with dplyr and data.table, the choice between Armadillo and Eigen As an example, the economiccomplexity package (Vargas Sepulveda 2020) uses base R depends on the user’s needs and preferences. For instance, Armadillo or Eigen can be ideal to work with a $000\times 1,000,$ 000 × 1 , 000 , 000 matrix but R can be more suitable for a $000\times 1,$ 000 × 1 , 000 matrix, and something similar applies to dplyr that is suitable for a 2-4 GB CSV files or SQL data but data.table is more suitable for 100 GB CSV datasets.
5 Cases where Armadillo and Eigen stand out
Vargas Sepulveda (2020) uses base R and the Matrix package to calculate the Balassa index and provides international trade data for 226 countries and 785 exported commodities.
Let $X\in\mathbb{R}^{C\times P}$ be a matrix with entries $x_{c,p}$ that represents the exports of country $c$ in product $p$, from this matrix the Balassa indices matrix is calculated as:
$B=([X\oslash(X\vec{1}_{P\times 1})]^{t}\oslash[X^{t}\vec{1}_{C\times 1}\oslash(\vec{1}_{C\times 1}^{t}X\vec{1}_{P\times 1})])^{t},$ (1)
[…]
This is the same as the Balassa index for country $c$ and product $p$:
$B_{cp}=\frac{x_{cp}}{\sum_{c}x_{cp}}/\frac{\sum_{p}x_{cp}}{\sum_{c}\sum_{p}x_{cp}}$ (2)
What is often used is to produce a zeroes and ones matrix $S$ defined as:
$s_{c,p}=\begin{cases}1&\text{ if }b_{cp}>1\cr 0&\text{ otherwise}\end{cases}$ (3)
(3) can be implemented in base R as:
balassa_r <- function(X) {
  B <- t(t(X / rowSums(X)) / (colSums(X) / sum(X)))
  B[B < 1] <- 0
  B[B >= 1] <- 1
  B
The C++ code using cpp11armadillo is:
#include <cpp11.hpp>
#include <cpp11armadillo.hpp>
using namespace cpp11;
using namespace arma;
[[cpp11::register]] doubles_matrix<> balassa_arma_(
  const doubles_matrix<>& x) {
  mat X = as_Mat(x);
  mat B = X.each_col() / sum(X, 1);
  B = B.each_row() / (sum(X, 0) / accu(X));
  B.elem(find(B < 1)).zeros();
  B.elem(find(B >= 1)).ones();
  return as_doubles_matrix(B);
The C++ code using cpp11eigen is:
#include <cpp11.hpp>
#include <cpp11eigen.hpp>
using namespace cpp11;
using namespace Eigen;
[[cpp11::register]] doubles_matrix<> balassa_eigen_(
  const doubles_matrix<>& x) {
  MatrixXd X = as_Matrix(x);
  MatrixXd B = X.array().rowwise() / X.rowwise().sum().array();
  B = B.array().colwise() / (X.colwise().sum().array() / X.sum());
  B = (B.array() < 1).select(0, B);
  B = (B.array() >= 1).select(1, B);
  return as_doubles_matrix(B);
If we use UN COMTRADE data for the year 2020 for 234 countries and 5,386 countries (United Nations 2023), we can observe that Armadillo and Eigen are around two times faster than base R at obtaining the Balassa matrix, and this includes the time to move the data between R and C++:
Table 4: Balassa indices
Operation Time (s) Rank
Balassa indices Eigen 0.013 1
Balassa indices Armadillo 0.014 2
Balassa indices R 0.026 3
The rest of the methods in Vargas Sepulveda (2020) involve recursion and eigenvalues computation, and these tasks were already covered in the ATT benchmark, meaning that the same speed gains can be expected as in the Balassa matrix.
6 Conclusion
Armadillo and Eigen can be highly expressive, these are flexible libraries once the user has learned the syntax, and these languages have data structures that do not exist in R that help to write efficient code. Eigen and cpp11eigen do not simplify the process of writing C++ code for R users but excels at computationally demanding applications. Armadillo and cpp11armadillo, on the other hand, provides a balance between speed and ease of use, and it is a good choice for users who need to write C++ code that is easier to modify and maintain.