### v2

# Welding R and C++: A Tale of Two Programming Languages

[…]
mm[#1]
Welding R and C++: A Tale of Two Programming Languages
Mauricio Vargas Sepúlveda (ORCID $0$-0003-1017-7574)
Department of Political Science, University of Toronto
Munk School of Global Affairs and Public Policy, University of Toronto
Corresponding author: m.sepulveda@mail.utoronto.ca
[…]
1 Abstract
This article compares cpp11armadillo and cpp11eigen, new R packages that integrate the powerful Armadillo and Eigen C++ libraries for linear algebra into the R programming environment. This article provides a detailed comparison between Armadillo and Eigen speed and syntax. The goal of these packages is to simplify a part of the process of solving bottlenecks by using C++ within R, these offer additional ease of integration for users who require high-performance linear algebra operations in their R workflows. This document aims to discuss the tradeoff between computational efficiency and accessibility.
2 Introduction
R is widely used by non-programmers (Wickham et al. 2019), and this article aims to introduce benchmarks in a non-technical yet formal manner for social scientists. Our goal is to provide a fair comparison between Eigen and Armadillo, being both highly efficient linear algebra libraries written in C++. We do it by using cpp11armadillo and cpp11eigen.
Armadillo is a C++ library designed for linear algebra, emphasizing a balance between performance and ease of use. C++ is highly efficient for computationally intensive tasks but lacks built-in data structures and functions for linear algebra operations. Armadillo fills this gap by providing an intuitive syntax similar to MATLAB (Sanderson and Curtin 2016).
Eigen emphasizes flexibility and speed, while Armadillo focuses on a balance between speed and easy of use.
RcppArmadillo, introduced in 2010, integrates R and Armadillo (Sanderson and Curtin 2016; Eddelbuettel and Sanderson 2014). RcppEigen, introduced in 2011, integrates Eigen with R through the Rcpp package introduced in 2008, enabling the use of C++ for performance-critical parts of R code. At the time of writing this document, 732 CRAN packages depend on RcppArmadillo and 238 on RcppEigen (Lee 2024), and therefore these are highly successful packages considering that the median number of reverse dependencies for CRAN packages is five and the distribution of dependencies situates these packages as top one percent.
cpp11armadillo and cpp11eigen are independent project that aim to simplify the integration of R and C++ by using cpp11, an R package introduced in 2020 that eases calling C++ functions from R, it is currently used as a dependency by 75 CRAN packages, and it is in the top one percent of CRAN packages (R Core Team 2024).
A distinctive characteristic of cpp11armadillo and cpp11eigen is the vendoring capability, meaning that it allows to copy its code into a project, making it a one-time dependency with a fixed and stable code until it is manually updated. This feature is useful in restricted environments such as servers and clusters where sometimes there are restrictions to software installation or the internet connections are limited for security reasons (Wickham et al. 2019; Vaughan, Hester, and François 2023).
cpp11armadillo and cpp11eigen are useful in cases where vectorization (e.g., applying an operation to a vector or matrix as a whole instead of looping over each element) is not possible or challenging. A detailed discussion and examples about why and when (and when not) rewriting R code in C++ is useful can be found in Burns (2011). We followed four design principles when developing these two packages as in: column oriented, package oriented, header-only, and vendoring capable. The details of the cpp11armadillo implementation, which is similar to cpp11eigen, can be found in Vargas Sepúlveda and Schneider Malamud (2024).
3 Syntax
One possibility is to start by creating minimal R packages with the provided templates. These templates provide a general case for a package that includes the necessary files to create a package that uses Armadillo or Eigen, including a generic Makevars file that can be adapted to link to specific numerical libraries such as Intel MKL or OpenBLAS.
remotes::install_github("pachadotdev/cpp11armadillo")
remotes::install_github("pachadotdev/cpp11eigen")
cpp11eigen::create_package("armadillobenchmark")
cpp11eigen::create_package("eigenbenchmark")
Comparing numerical libraries requires to write equivalent codes. For instance, in R we use apply() while its C++ equivalent is a for loop, and this allows a fair comparison between the two libraries. However, R has heavily optimized functions that also verify the input data, such as lm() and glm(), which do not have a direct equivalent in Armadillo or Eigen, and for a fair comparison the options are to write a simplified function for the linear model in R or to write a more complex function in C++.
The ATT benchmark, is a set of functions that can be rewritten using Armadillo and Eigen with relative ease, and test has the advantage of being well-known and widely used in the R community.
The first test in the ATT benchmark is the creation, transposition and deformation of an $N\times N$ matrix ($500\times 2,$ 500 × 2 , 500 in the original test). The R and Armadillo codes for this operation are:
# R
matrix_calculation_01_r <- function(n) {
  a <- matrix(rnorm(n * n) / 10, ncol = n, nrow = n)
  b <- t(a)
  dim(b) <- c(n / 2, n * 2)
  a <- t(b)
  return(0L)
// C++
#include <cpp11.hpp>
#include <cpp11armadillo.hpp>
using namespace arma;
using namespace cpp11;
[[cpp11::register]] int matrix_calculation_01_arma_(const int& n) {
  mat a = randn<mat>(n,n) / 10;
  mat b = a.t();
  b.reshape(n/2, n*2);
  a = b.t();
  return 0;
The Eigen code requires to create a function to draw random numbers from a normal distribution because it only provides a built-in function for the uniform distribution:
#include <cpp11.hpp>
#include <cpp11eigen.hpp>
#include <random>
using namespace Eigen;
using namespace cpp11;
std::mt19937& random_normal() {
  static std::random_device rd;
  static std::mt19937 gen(rd());
  return gen;
[[cpp11::register]] int matrix_calculation_01_eigen_(const int& n) {
  std::normal_distribution<double> d(0, 1);
  MatrixXd a = MatrixXd::NullaryExpr(n, n, [&]() {
    return d(random_normal());
  }) / 10;
  // for the uniform distribution this is just
  // MatrixXd a = MatrixXd::Random(n, n) / 10;
  MatrixXd b = a.transpose();
  b.resize(n / 2, n * 2);
  return 0;
4 Benchmarks without data transfer
The functions in the previous section to do not move data between R and C++, this is intentional to focus on the performance of the linear algebra libraries and not adding overhead from data transfer in the benchmarks. Each function creates a matrix and conducts equivalent operations on it. The returned value is zero in R and C++ in case that the functions run without errors.
We decided to run the benchmarks on a local machine and on a cluster to provide a comparison between the two environments and test how the benchmarks change when the input data is increased.
4.1 Local benchmarks
The local benchmarks were conducted on a ThinkPad X1 Carbon Gen 9 with the following specifications:
  • Processor: Intel Core i7-1185G7 with eight cores
  • Memory: 16 GB LPDDR4Xx-4266
  • Operating System: Pop!_OS 22.04 based on Ubuntu 22.04
  • R Version: 4.4.1
  • BLAS Library: OpenBLAS 0.3.20
The median times for the adapted and comparable implementations of the ATT benchmarks are as follows:
Table 1: Matrix calculation
Operation Median time (s) Rank
$400\times 2,$ 400 × 2 , 400 $\text{matrix}^{1,000}$ Armadillo $0.188$ 1
$400\times 2,$ 400 × 2 , 400 $\text{matrix}^{1,000}$ Eigen $0.301$ 2
$400\times 2,$ 400 × 2 , 400 $\text{matrix}^{1,000}$ R $0.325$ 3
$800\times 2,$ 800 × 2 , 800 cross-product matrix - Armadillo $0.398$ 1
$800\times 2,$ 800 × 2 , 800 cross-product matrix - R $0.444$ 2
$800\times 2,$ 800 × 2 , 800 cross-product matrix - Eigen $1.151$ 3
Creation and modification of a $500\times 2,$ 500 × 2 , 500 matrix - Armadillo $0.204$ 1
Creation and modification of a $500\times 2,$ 500 × 2 , 500 matrix - Eigen $0.232$ 2
Creation and modification of a $500\times 2,$ 500 × 2 , 500 matrix - R $0.294$ 3
Linear regression over a $000\times 3,$ 000 × 3 , 000 matrix - Armadillo $0.459$ 1
Linear regression over a $000\times 3,$ 000 × 3 , 000 matrix - R $5.303$ 2
Linear regression over a $000\times 3,$ 000 × 3 , 000 matrix - Eigen $8.809$ 3
Sorting of 7,000,000 values - Armadillo $0.663$ 1
Sorting of 7,000,000 values - Eigen $0.691$ 2
Sorting of 7,000,000 values - R $0.759$ 3
Table 2: Matrix functions
Operation Median time (s) Rank
Cholesky decomposition of a $000\times 3,$ 000 × 3 , 000 matrix - Armadillo $0.608$ 1
Cholesky decomposition of a $000\times 3,$ 000 × 3 , 000 matrix - R $0.709$ 2
Cholesky decomposition of a $000\times 3,$ 000 × 3 , 000 matrix - Eigen $2.902$ 3
Determinant of a $500\times 2,$ 500 × 2 , 500 matrix - Armadillo $0.293$ 1
Determinant of a $500\times 2,$ 500 × 2 , 500 matrix - R $0.303$ 2
Determinant of a $500\times 2,$ 500 × 2 , 500 matrix - Eigen $0.562$ 3
Eigenvalues of a $640\times 640$ matrix - Armadillo $0.367$ 1
Eigenvalues of a $640\times 640$ matrix - R $0.369$ 2
Eigenvalues of a $640\times 640$ matrix - Eigen $1.629$ 3
Fast Fourier Transform over 2,400,000 values - Eigen $0.14$ 1
Fast Fourier Transform over 2,400,000 values - R $0.23$ 2
Fast Fourier Transform over 2,400,000 values - Armadillo $0.294$ 3
Inverse of a $600\times 1,$ 600 × 1 , 600 matrix - Armadillo $0.312$ 1
Inverse of a $600\times 1,$ 600 × 1 , 600 matrix - R $0.324$ 2
Inverse of a $600\times 1,$ 600 × 1 , 600 matrix - Eigen $0.758$ 3
Table 3: Programmation
Operation Median time (s) Rank
3,500,000 Fibonacci numbers calculation - Eigen $1.4\times 10^{-1}$ 1
3,500,000 Fibonacci numbers calculation - Armadillo $1.7\times 10^{-1}$ 2
3,500,000 Fibonacci numbers calculation - R $1.7\times 10^{-1}$ 3
Creation of a $000\times 3,$ 000 × 3 , 000 Hilbert matrix - Eigen $4.6\times 10^{-6}$ 1
Creation of a $000\times 3,$ 000 × 3 , 000 Hilbert matrix - Armadillo $5.9\times 10^{-2}$ 2
Creation of a $000\times 3,$ 000 × 3 , 000 Hilbert matrix - R $1.5\times 10^{-1}$ 3
Creation of a $500\times 500$ Toeplitz matrix - Eigen $7.9\times 10^{-7}$ 1
Creation of a $500\times 500$ Toeplitz matrix - Armadillo $4\times 10^{-4}$ 2
Creation of a $500\times 500$ Toeplitz matrix - R $2.6\times 10^{-3}$ 3
Escoufier’s method on a $45\times 45$ matrix - Armadillo $2.4\times 10^{-2}$ 1
Escoufier’s method on a $45\times 45$ matrix - Eigen $3.2\times 10^{-2}$ 2
Escoufier’s method on a $45\times 45$ matrix - R $1.4\times 10^{-1}$ 3
Grand common divisors of 400,000 pairs - Eigen $2.1\times 10^{-2}$ 1
Grand common divisors of 400,000 pairs - Armadillo $2.3\times 10^{-2}$ 2
Grand common divisors of 400,000 pairs - R $1.884$ 3
The results reveal that Armadillo leads in most of the benchmarks, but Eigen is particularly faster in some tests such as the Fast Fourier Transform. R is the second or third in all benchmarks, but it is important to note that R comes with an additional advantage in terms of simplified syntax and the ability to run the code without compiling it.
These tests are not exhaustive, and we must be cautious when interpreting the results. The ATT benchmark is a good starting point, but it does not cover mundane tasks such as data manipulation, and it is important to consider the tradeoff between computational efficiency and ease of use.
4.2 Cluster benchmarks
The cluster benchmarks were conducted on one cluster node of the Niagara supercomputer maintained by the Digital Research Alliance of Canada, which has the following specifications:
  • Processor: 2 sockets with 20 Intel Skylake cores (2.4GHz, AVX512), for a total of 40 cores per node
  • Memory: 202 GB
  • Operating System: CentOS 7
  • R Version: 4.2.2
  • BLAS Library: Intel MKL 2019.4.243
The median times for the adapted and comparable implementations of the ATT benchmarks are as follows:
Table 4: Matrix calculation
Operation Median time (s) Rank
$000\times 12,$ 000 × 12 , 000 $\text{matrix}^{1,000}$ Eigen $0.564$ 1
$000\times 12,$ 000 × 12 , 000 $\text{matrix}^{1,000}$ Armadillo $0.763$ 2
$000\times 12,$ 000 × 12 , 000 $\text{matrix}^{1,000}$ R $0.988$ 3
$000\times 14,$ 000 × 14 , 000 cross-product matrix - Armadillo $0.47$ 1
$000\times 14,$ 000 × 14 , 000 cross-product matrix - R $0.625$ 2
$000\times 14,$ 000 × 14 , 000 cross-product matrix - Eigen $1.322$ 3
Creation and modification of a $500\times 12,$ 500 × 12 , 500 matrix - Armadillo $0.321$ 1
Creation and modification of a $500\times 12,$ 500 × 12 , 500 matrix - Eigen $0.351$ 2
Creation and modification of a $500\times 12,$ 500 × 12 , 500 matrix - R $0.594$ 3
Linear regression over a $000\times 15,$ 000 × 15 , 000 matrix - Armadillo $0.616$ 1
Linear regression over a $000\times 15,$ 000 × 15 , 000 matrix - R $8.084$ 2
Linear regression over a $000\times 15,$ 000 × 15 , 000 matrix - Eigen $9.604$ 3
Sorting of 35,000,000 values - Armadillo $0411$
Sorting of 35,000,000 values - Eigen $1.067$ 2
Sorting of 35,000,000 values - R $1.32$ 3
Table 5: Matrix functions
Operation Median time (s) Rank
Cholesky decomposition of a $000\times 15,$ 000 × 15 , 000 matrix - Armadillo $0.585$ 1
Cholesky decomposition of a $000\times 15,$ 000 × 15 , 000 matrix - R $0.819$ 2
Cholesky decomposition of a $000\times 15,$ 000 × 15 , 000 matrix - Eigen $1.927$ 3
Determinant of a $500\times 12,$ 500 × 12 , 500 matrix - Armadillo $0.387$ 1
Determinant of a $500\times 12,$ 500 × 12 , 500 matrix - R $0.503$ 2
Determinant of a $500\times 12,$ 500 × 12 , 500 matrix - Eigen $0.635$ 3
Eigenvalues of a $200\times 3,$ 200 × 3 , 200 matrix - Armadillo $0.429$ 1
Eigenvalues of a $200\times 3,$ 200 × 3 , 200 matrix - R $0.464$ 2
Eigenvalues of a $200\times 3,$ 200 × 3 , 200 matrix - Eigen $2.602$ 3
Fast Fourier Transform over 12,000,000 values - Eigen $0.301$ 1
Fast Fourier Transform over 12,000,000 values - R $0.362$ 2
Fast Fourier Transform over 12,000,000 values - Armadillo $0.553$ 3
Inverse of a $000\times 8,$ 000 × 8 , 000 matrix - Armadillo $0.19$ 1
Inverse of a $000\times 8,$ 000 × 8 , 000 matrix - R $0.234$ 2
Inverse of a $000\times 8,$ 000 × 8 , 000 matrix - Eigen $0.448$ 3
Table 6: Programmation
Operation Median time (s) Rank
17,500,000 Fibonacci numbers calculation - Armadillo $5.9\times 10^{-1}$ 1
17,500,000 Fibonacci numbers calculation - Eigen $5.9\times 10^{-1}$ 2
17,500,000 Fibonacci numbers calculation - R $7.1\times 10^{-1}$ 3
Creation of a $000\times 15,$ 000 × 15 , 000 Hilbert matrix - Eigen $3.1\times 10^{-2}$ 1
Creation of a $000\times 15,$ 000 × 15 , 000 Hilbert matrix - Armadillo $3.7\times 10^{-2}$ 2
Creation of a $000\times 15,$ 000 × 15 , 000 Hilbert matrix - R $2.4\times 10^{-1}$ 3
Creation of a $500\times 2,$ 500 × 2 , 500 Toeplitz matrix - Eigen $3.1\times 10^{-4}$ 1
Creation of a $500\times 2,$ 500 × 2 , 500 Toeplitz matrix - Armadillo $5.0\times 10^{-4}$ 2
Creation of a $500\times 2,$ 500 × 2 , 500 Toeplitz matrix - R $4.1\times 10^{-3}$ 3
Escoufier’s method on a $225\times 225$ matrix - Eigen $3.4\times 10^{-2}$ 1
Escoufier’s method on a $225\times 225$ matrix - Armadillo $4.8\times 10^{-2}$ 2
Escoufier’s method on a $225\times 225$ matrix - R $3.0\times 10^{-1}$ 3
Grand common divisors of 2,000,000 pairs - Armadillo $3.8\times 10^{-2}$ 1
Grand common divisors of 2,000,000 pairs - Eigen $4.0\times 10^{-2}$ 2
Grand common divisors of 2,000,000 pairs - R $2.793$ 3
Repeating the same after multiplicating the number of rows and columns by five leads to the following results:
Table 7: Matrix calculation
Operation Median time (s) Rank
$000\times 12,$ 000 × 12 , 000 $\text{matrix}^{1,000}$ Eigen $3.9741$
$000\times 12,$ 000 × 12 , 000 $\text{matrix}^{1,000}$ Armadillo $18.988$ 2
$000\times 12,$ 000 × 12 , 000 $\text{matrix}^{1,000}$ R $24.48$ 3
$000\times 14,$ 000 × 14 , 000 cross-product matrix - Armadillo $3.781$
$000\times 14,$ 000 × 14 , 000 cross-product matrix - R $17.7$ 2
$000\times 14,$ 000 × 14 , 000 cross-product matrix - Eigen $127.95$ 3
Creation and modification of a $500\times 12,$ 500 × 12 , 500 matrix - Armadillo $8.037$ 1
Creation and modification of a $500\times 12,$ 500 × 12 , 500 matrix - Eigen $8.797$ 2
Creation and modification of a $500\times 12,$ 500 × 12 , 500 matrix - R $13.98$ 3
Linear regression over a $000\times 15,$ 000 × 15 , 000 matrix - Armadillo $6.6341$
Linear regression over a $000\times 15,$ 000 × 15 , 000 matrix - R $1265$ 2
Linear regression over a $000\times 15,$ 000 × 15 , 000 matrix - Eigen $1517.8$ 3
Sorting of 35,000,000 values - Armadillo $5.556$ 1
Sorting of 35,000,000 values - Eigen $5.708$ 2
Sorting of 35,000,000 values - R $6.952$ 3
Table 8: Matrix functions
Operation Median time (s) Rank
Cholesky decomposition of a $000\times 15,$ 000 × 15 , 000 matrix - Armadillo $6.8211$
Cholesky decomposition of a $000\times 15,$ 000 × 15 , 000 matrix - R $1.0272$
Cholesky decomposition of a $000\times 15,$ 000 × 15 , 000 matrix - Eigen $184.11$ 3
Determinant of a $500\times 12,$ 500 × 12 , 500 matrix - Armadillo $0.0041$
Determinant of a $500\times 12,$ 500 × 12 , 500 matrix - R $12.981$ 2
Determinant of a $500\times 12,$ 500 × 12 , 500 matrix - Eigen $40.249$ 3
Eigenvalues of a $200\times 3,$ 200 × 3 , 200 matrix - Armadillo $9.11$ 1
Eigenvalues of a $200\times 3,$ 200 × 3 , 200 matrix - R $9.445$ 2
Eigenvalues of a $200\times 3,$ 200 × 3 , 200 matrix - Eigen $519.36$ 3
Fast Fourier Transform over 12,000,000 values - Eigen $751$
Fast Fourier Transform over 12,000,000 values - R $3332$
Fast Fourier Transform over 12,000,000 values - Armadillo $5293$
Inverse of a $000\times 8,$ 000 × 8 , 000 matrix - Armadillo $5.077$ 1
Inverse of a $000\times 8,$ 000 × 8 , 000 matrix - R $6.846$ 2
Inverse of a $000\times 8,$ 000 × 8 , 000 matrix - Eigen $6.6933$
Table 9: Programmation
Operation Median time (s) Rank
17,500,000 Fibonacci numbers calculation - Armadillo $2.893$ 1
17,500,000 Fibonacci numbers calculation - Eigen $972$
17,500,000 Fibonacci numbers calculation - R $5763$
Creation of a $000\times 15,$ 000 × 15 , 000 Hilbert matrix - Eigen $9.2\times 10^{-1}$ 1
Creation of a $000\times 15,$ 000 × 15 , 000 Hilbert matrix - Armadillo $1.062$ 2
Creation of a $000\times 15,$ 000 × 15 , 000 Hilbert matrix - R $4.428$ 3
Creation of a $500\times 2,$ 500 × 2 , 500 Toeplitz matrix - Eigen $1.8\times 10^{-2}$ 1
Creation of a $500\times 2,$ 500 × 2 , 500 Toeplitz matrix - Armadillo $2.9\times 10^{-2}$ 2
Creation of a $500\times 2,$ 500 × 2 , 500 Toeplitz matrix - R $1.6\times 10^{-1}$ 3
Escoufier’s method on a $225\times 225$ matrix - Armadillo $25.369$ 1
Escoufier’s method on a $225\times 225$ matrix - Eigen $38.911$ 2
Escoufier’s method on a $225\times 225$ matrix - R $403.24$ 3
Grand common divisors of 2,000,000 pairs - Armadillo $1.9\times 10^{-1}$ 1
Grand common divisors of 2,000,000 pairs - Eigen $2.0\times 10^{-1}$ 2
Grand common divisors of 2,000,000 pairs - R $15.715$ 3
The results are consistent with the local benchmarks, and Armadillo leads in most of the tests. The benchmarks are also consistent with the time complexity of the algorithms, meaning that doubling the size of the matrix does not double the time to run the function unless the function has a time complexity of $O(n)$.
5 Benchmarks with data transfer
Psarras, Barthels, and Bientinesi (2022) provides different benchmarks for the Linear Algebra Mapping Problem. We have adapted their benchmarks to solve linear systems in order to extrapolate their findings to a larger input data.
We repeated the experiment consisting in solving a linear system of equations $AX=B$ with $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times m}$ for $n=30,000$ and $m=1,000$. Because only cpp11armadillo has available methods to pass sparse matrices between R and C++, we created dense matrices to also include cpp11eigen in the comparison, and this also adds additional stress to the tests.
A dense matrix with double precision entries $A\in\mathbb{R}^{30,000\times 30,000}$ uses 6.7 GB of RAM and can be created and operated in the Niagara cluster without problems. The benchmark repeats the same task in two different ways, in a naive way by directly computing $A^{-1}B$ and in a smart way by using solve(A,B). The solution always exists because the data is created by first defining random matrices $A$ and $X$ and then computing $B=AX$, the benchmarks use $A$ and $B$ as inputs to solve for $X$. For the cases where $A^{-1}$ is not directly obtained, R and Armadillo use the LU decomposition by inspecting the matrix structure while in Eigen this has to be specified, and this is the correct approach provided that $A$ was created with the rnorm() function in R without guarantee of symmetry nor it results in an overdetermined system.
The benchmark results are as follows:
Table 10: Solving linear systems
Operation Median time (s) Rank
Naive solution - Armadillo 18.358 1
Naive solution - Base R 62.138 2
Naive solution - Eigen 1899.225 3
Smart solution - Base R 25.494 1
Smart solution - Armadillo 29.394 2
Smart solution - Eigen 467.490 3
Armadillo leads in performance for the naive solution. This means that, even after considering the overhead of data transfer, Armadillo excels at computationally intensive tasks that involve repeated operations as it is the case of repeating row and column multiplication to obtain $X$, and this benchmark does not cover additional data structures in C++ that do not exist in R and that provide additional flexibility.
R leads in performance for the smart solution not because of the data transfer overhead, but because R internally uses LAPACK and BLAS libraries that are highly optimized for linear algebra operations. In the particular case of the Niagara cluster, R was compiled against the Intel MKL library, which is highly optimized for Intel processors and benefits from the specific processor instructions set. Armadillo and Eigen in this case also use the Intel MKL library, and for a benchmark with a smaller size input data (e.g. $100\times 100$)the overhead of data transfer and using 40 cores would be higher than the speed gains.
6 Similarities between C++ libraries and R packages
The syntax and speed differences when using Armadillo or Eigen in C++ posit a similar case to the tradeoff between using dplyr and data.table (Wickham et al. 2019; Barrett et al. 2024). dplyr is easier to use but data.table is faster. dplyr was not designed to be fast but data.table was not designed to be easy to use. For instance, the code to obtain the grouped means by number of cylinders in the mtcars dataset is:
# dplyr
mtcars %>%
  group_by(cyl) %>%
  summarise_all(mean)
# data.table
as.data.table(mtcars)[, lapply(.SD, mean), by = cyl]
The local benchmark for the grouped means reveals that dplyr has a median time of 2.7 ms and data.table has a median time of 600 $\mu$s, meaning that dplyr is four times slower than data.table at this task. The syntax of dplyr is easier to understand for non-programmers, but data.table can be equally expressive for users who are familiar with its syntax.
The tests for Armadillo and Eigen reveal that, for repeated and computationally intensive tasks, rewriting R code in C++ can lead to significant performance improvements, but it comes at the cost of learning a new syntax.
As with dplyr and data.table, the choice between Armadillo and Eigen depends on the user’s needs and preferences. For instance, Armadillo or Eigen can be ideal to work with a $000\times 1,000,$ 000 × 1 , 000 , 000 matrix but R can be more suitable for a $000\times 1,$ 000 × 1 , 000 matrix, and something similar applies to dplyr that is suitable for a 2-4 GB CSV files or SQL data but data.table is more suitable for large datasets (e.g., 100 GB CSV files).
7 Cases where Armadillo and Eigen stand out
Using Armadillo or Eigen can be particularly useful for functions that involve nested loops and recursion. If we are going to repeatedly use a function that requires nested loops or multiple linear algebra operations, it may be worth rewriting it in C++ using Armadillo or Eigen instead of using base R. In such cases, the time incurred in learning the syntax and obtaining the correct function is an investment that pays off in long run time savings.
For instance, Vargas Sepulveda (2020) uses base R and the Matrix package to calculate the Balassa index and provides international trade data for 226 countries and 785 exported commodities. A matrix of $226\times 785$ does not pose a problem for base R, nor it counts as big data, but it shows large speed gains when using Armadillo or Eigen.
Let $X\in\mathbb{R}^{C\times P}$ be a matrix with entries $x_{c,p}$ that represents the exports of country $c$ in product $p$, from this matrix the Balassa indices matrix is calculated as:
$B=([X\oslash(X\vec{1}_{P\times 1})]^{t}\oslash[X^{t}\vec{1}_{C\times 1}\oslash(\vec{1}_{C\times 1}^{t}X\vec{1}_{P\times 1})])^{t},$ (1)
[…]
This is the same as the Balassa index for country $c$ and product $p$:
$B_{cp}=\frac{x_{cp}}{\sum_{c}x_{cp}}/\frac{\sum_{p}x_{cp}}{\sum_{c}\sum_{p}x_{cp}}.$ (2)
$B$ is often used to produce a zeroes and ones matrix $S$ defined as:
$s_{c,p}=\begin{cases}1&\text{ if }b_{cp}>1\cr 0&\text{ otherwise}\end{cases},$ (3)
where a value of one indicates that country $c$ has a revealed comparative advantage in product $p$ and zero otherwise.
(3) can be implemented in base R as:
balassa_r <- function(X) {
  B <- t(t(X / rowSums(X)) / (colSums(X) / sum(X)))
  B[B < 1] <- 0
  B[B >= 1] <- 1
  B
The C++ code using cpp11armadillo is:
#include <cpp11.hpp>
#include <cpp11armadillo.hpp>
using namespace cpp11;
using namespace arma;
[[cpp11::register]] doubles_matrix<> balassa_arma_(
  const doubles_matrix<>& x) {
  mat X = as_Mat(x);
  mat B = X.each_col() / sum(X, 1);
  B = B.each_row() / (sum(X, 0) / accu(X));
  B.elem(find(B < 1)).zeros();
  B.elem(find(B >= 1)).ones();
  return as_doubles_matrix(B);
The C++ code using cpp11eigen is:
#include <cpp11.hpp>
#include <cpp11eigen.hpp>
using namespace cpp11;
using namespace Eigen;
[[cpp11::register]] doubles_matrix<> balassa_eigen_(
  const doubles_matrix<>& x) {
  MatrixXd X = as_Matrix(x);
  MatrixXd B = X.array().rowwise() / X.rowwise().sum().array();
  B = B.array().colwise() / (X.colwise().sum().array() / X.sum());
  B = (B.array() < 1).select(0, B);
  B = (B.array() >= 1).select(1, B);
  return as_doubles_matrix(B);
If we use UN COMTRADE data for the year 2020 for 234 countries and 5,386 countries in the finest granularity level data from United Nations (2023), we can observe that Armadillo and Eigen are around two times faster than base R at obtaining the Balassa matrix, and this includes the time to move the data between R and C++:
Table 11: Balassa indices
Operation Median time (s) Rank
Balassa indices Eigen 0.013 1
Balassa indices Armadillo 0.014 2
Balassa indices R 0.026 3
The rest of the methods in Vargas Sepulveda (2020) involve recursion and eigenvalues computation, and these tasks were already covered in the ATT benchmark, meaning that the same speed gains can be expected as in the Balassa matrix.
8 Considerations
Being fair, R was not designed to be fast, and it is not a low-level language like C++. R is a high-level language that some users consider easy to learn and use, and it is particularly useful for data manipulation and visualization. R has a large number of packages that can be used to conduct a wide range of statistical analyses, and it is particularly useful for non-programmers.
Armadillo and Eigen are not designed to be easy to use, and they involve a learning curve to use them effectively. These libraries are particularly useful for computationally intensive tasks that involve nested loops and recursion. These languages have speed advantages over R for two main reasons, that the time it takes for each step in a loop is shorter in C++ than in R, and that these libraries provide efficient data structures for vectors and matrices besides providing internal methods that combine operations to increase speed and reduce memory usage.
The choice between R and C++ depends on the user’s needs and preferences. In a way, comparing R and C++ is like comparing a Vespa and a Ducati motorcycle, both are motorcycles but they are designed for different purposes and excel in different areas. For instance, a Vespa is ideal for city commuting and a Ducati is ideal for racing, and the same applies to R and C++.
Using C++ for tasks such as data manipulation and visualization is like using a Ford F-150 to go to the convenience store, it is feasible but it is not the most efficient way to do it and the additional fuel consumption compared to a lighter vehicle ressembles the extra time required to compile and corrrect the code. In the same way, using R for computationally intensive tasks is like using a Mini Cooper to load bricks and construction materials, it is feasible but you cannot expect the same experience (e.g., speed and comfort) as using a Ford F-150 for the same task.
Writing proper C++ code requires a good understanding of the syntax and reading the respective documentation that each library provides. For instance, without reading the Armadillo documentation, it is very tempting to transpose a matrix in the following way:
mat bad_transposition_(const int& n) {
  mat a = randn<mat>(n, n) / 10;
  mat b(a.n_cols, a.n_rows);
  for (int i = 0; i < a.n_rows; i++) {
    for (int j = 0; j < a.n_cols; j++) {
      b(i, j) = a(j, i);
  return b;
Or, even worse, trying to make it faster by using OMP, which would be faster but not efficient in terms of syntax and available alternatives:
mat less_bad_transposition_(const int& n) {
  mat a = randn<mat>(n, n) / 10;
  mat b(a.n_cols, a.n_rows);
  #ifdef _OPENMP
  #pragma omp parallel for collapse(2) schedule(static)
  #endif
  for (int i = 0; i < a.n_rows; i++) {
    for (int j = 0; j < a.n_cols; j++) {
      b(i, j) = a(j, i);
  return b;
The correct way to transpose a matrix in Armadillo, such that its internals apply low-level optimizations, is to use the t() function, and it also saves time and typing:
mat good_transposition_(const int& n) {
  mat a = randn<mat>(n, n) / 10;
  return a.t();
9 Conclusion
Armadillo and Eigen can be highly expressive, these are flexible libraries once the user has learned the syntax, and these languages have data structures that do not exist in R that help to write efficient code. Eigen and cpp11eigen do not simplify the process of writing C++ code for R users but excels at computationally demanding applications. Armadillo and cpp11armadillo, on the other hand, provides a balance between speed and ease of use, and it is a good choice for users who need to write C++ code that is easier to modify and maintain.
10 Acknowledgements
I appreciate that Paolo Bientinesi from Umea University read a previous version of this document on arXiv and referred me to Psarras, Barthels, and Bientinesi (2022).